OpenStudy (anonymous):

Einstein's formula (E=mc^2) tells us how much energy is emitted when mass is annihilated, where E is energy (in joules), m is mass and c is the speed of light (3 x 10^8 m/s). A penny has a mass of 2.35 g. How much Energy would be produced if it was converted into energy? The final answer must be in kWh

OpenStudy (jamesj):

E = mc^2 = (2.35 )( 3 x 10^8)^2 J Now, Power = Energy/Time 1 Watt = 1 Joule/1 sec = 1 J/s hence 1 Watt.hour = 1 Wh = 60^2 J/ = 3,600 J 1 kilo Watt.hour = 1 kWh = 1000 x 3,600 J = 3,600,000 J So once you've found the energy above, divide it by 3,600,000 to express it in kWh.

OpenStudy (jamesj):

Make sense?

OpenStudy (anonymous):

yep. tnx ;)

OpenStudy (anonymous):

you should convert g to kg into so, 2.35 g=0.00235 kg

OpenStudy (jamesj):

That's correct. So the energy is E = mc^2 = (0.00235 )( 3 x 10^8)^2 J Thanks for the catch.

OpenStudy (anonymous):

E=2*10^14 J 1J=3*10-7 kWh so \[E \approx 6*10^{7} kWh \]

OpenStudy (anonymous):

1J=3*10^-7 kWh